OpenStudy (anonymous):

An earthquake registered 7.4 on the Richter scale. If the reference intensity of this quake was 2.0 × 10^11, what was its intensity?

OpenStudy (anonymous):

so 2.0x10^11 corrrespond to 1 in the scale?

OpenStudy (anonymous):

Mmhm

OpenStudy (anonymous):

i think it is so that is a proportion problem

OpenStudy (anonymous):

if 2.0x10^11 corrrespond to 1 in the scale? x correspond to 7.4 \[\frac{ 2.0x10^(11)}{ x } =\frac{ 1 }{ 7.4 }\]